Indicate in the row of electrolytes acids, hydroxides and salts: CdCl2, Mn(OH)2, Fe(NO3)2, H2SiO3, KOH, C2H5COOH, Ca(OH)2, HCNS, NiHPO4. Select weak electrolytes and write the equations of step dissociation and expressions of step dissociation constants. Write the general equations of dissociation and expressions of the general dissociation constants.
1
Expert's answer
2020-12-14T14:35:12-0500

Acids - H2SiO3 , C2H5COOH , HCNS

Hydroxides -  Mn(OH)2 , KOH, Ca(OH)2

Salts - CdCl2 , Fe(NO3 )2 , NiHPO4


C2H5COOH is a weak acid.


C2H5COOH "\to" C2H5COO- + H+

Dissociation constant = "\frac{[C2H5COO^-][H^+]}{[C2H5COOH]}" All the conc. are at taken at equilibrium.


H+ + C2H5COOH"\to" C2H5COOH H+

Dissociation constant = "\frac{[C2H5COOH H^+]}{[C2H5COOH][H^+]}"All the conc. are at taken at equilibrium.


General equation : HA"\to" H+ + A_

Dissociation constant = "\frac{[H+][A^-]}{[HA]}" All the conc. are at taken at equilibrium.